OpenStudy (anonymous):

If f(x) = 2x2 + 3x and g(x) = x – 2, what is (f + g)(2)?

OpenStudy (michele_laino):

by definition, we can write this: \[\left( {f + g} \right)\left( x \right) = f\left( x \right) + g\left( x \right) = 2{x^2} + 3x + x - 2 = ...\] please complete

OpenStudy (anonymous):

2x squared + 4x-2??

OpenStudy (michele_laino):

ok! correct!

OpenStudy (anonymous):

i have some answer choices

OpenStudy (michele_laino):

now, please replace x with 2, what do you get? \[\begin{gathered} \left( {f + g} \right)\left( x \right) = 2{x^2} + 4x - 2 \hfill \\ \hfill \\ \left( {f + g} \right)\left( 2 \right) = 2 \times {2^2} + 4 \times 2 - 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

16 14 12 10 8 are the choices

OpenStudy (anonymous):

okay let me solve

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

is 12 correct?

OpenStudy (michele_laino):

hint: \[\left( {f + g} \right)\left( 2 \right) = 2 \times {2^2} + 4 \times 2 - 2 = 2 \times 4 + 8 - 2 = ...?\]

OpenStudy (anonymous):

2??

OpenStudy (michele_laino):

I got this: \[\begin{gathered} \left( {f + g} \right)\left( 2 \right) = 2 \times {2^2} + 4 \times 2 - 2 = \hfill \\ \hfill \\ = \left( {2 \times 4} \right) + 8 - 2 = 8 + 8 - 2 = 16 - 2 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

14???

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

thank you soooo much

OpenStudy (michele_laino):

:)